You are a farmer about to harvest the crop. To describe the uncertainty in the size of the harvest, you feel that it may be described as normal distribution with a mean of 80,000 bushels with a standard deviation  of 2500 bushels. Find the probability that your harvest will exceed 84,400 bushels.


1
Expert's answer
2021-02-24T12:20:38-0500

Solution:

Given, mean, "\mu=80,000"

And standard deviation, "\sigma=2500"

Let X denotes the random variable such that "X\sim N(80000,\ 2500)"

We need to find "P(X>84400)" .

"P(X>84400)=1-P(X\le84400)"

"=1-P(z\le \dfrac{84400-80000}{2500})" ["\because z=\dfrac{X-\mu}{\sigma}" ]

"=1-P(z\le \dfrac{4400}{2500})=1-P(z\le 1.76)"

"=1-0.96080=0.0392"

Thus, the required probability is 0.0392